OpenStudy (anonymous):

find the equation of the line thorugh the given pair of points in standard form using only intergers (5,2)and (-1,6)

OpenStudy (anonymous):

hellpp

zepdrix (zepdrix):

whut 0_o math?

zepdrix (zepdrix):

Sooo uhhh, this line you speak of... it has some points..

OpenStudy (anonymous):

y=mx+b

zepdrix (zepdrix):

Standard form? Mmm that's no fun.

zepdrix (zepdrix):

Slope-intercept form? Ok good good, we can work with that :)

OpenStudy (anonymous):

m=-1/3

OpenStudy (anonymous):

-2/3

zepdrix (zepdrix):

\[\Large\rm (5,2), \qquad (-1,6)\] \(\Large\rm m=\frac{6-2}{-1-5}=-\frac{2}{3}\) Mmm ok good, that looks better :)

OpenStudy (anonymous):

2=-2/3(5)+b

zepdrix (zepdrix):

Mmmmm k, and b isssss? :O

OpenStudy (anonymous):

16/3

OpenStudy (anonymous):

??

zepdrix (zepdrix):

Ok great! Problem is almost done!

zepdrix (zepdrix):

\[\Large\rm y=-\frac{2}{3}x+\frac{16}{3}\]Now to put it into `standard form`. Do you remember what that looks like?

OpenStudy (anonymous):

2/3x+y=16/3 but wait i got that ... thats not one of my multiple choices lol

zepdrix (zepdrix):

Ok that's a good step. The final step would be to get rid of the fractions. We don't want fractions in standard form.

OpenStudy (anonymous):

my choices are a) 4x+6y=32 b) 4x-6y=32 c)6x+6y=32 d) 4c+6y=-32

OpenStudy (anonymous):

and ok

OpenStudy (anonymous):

but how

zepdrix (zepdrix):

Multiply both sides by the `Least Common Multiple` of your denominators. Since we only have 3's in the denominator, it makes it a bit easier. We'll multiply both sides by 3. Oops oops oops. Someone was being stupid when they made your multiple choice options, not your fault. So instead we'll multiply both sides by 6.

OpenStudy (anonymous):

k

zepdrix (zepdrix):

\[\Large\rm 3y=-3\frac{2}{3}x+3\frac{16}{3}\]Do you see the nice cancellations?

zepdrix (zepdrix):

Woops I did 3 when I told you to do 6 lolol

OpenStudy (anonymous):

its a

OpenStudy (anonymous):

:)))))

zepdrix (zepdrix):

Yay good job \c:/

OpenStudy (anonymous):

i haave one more question im lost on

zepdrix (zepdrix):

What'chu got? :U

OpenStudy (anonymous):

carl participates in a bowling tournament. after completing 5/8 of the game, his average score per game is 200. what must his average score be in the remaining games in order for his overall game to be 226.25? round your answer to the neitest point.

zepdrix (zepdrix):

Mmmm I've always hated these types of problems lol. Here is how I'm interpreting it. 5/8 of the 200 points, plus 3/8 of some unknown amount should add up to 8/8 of 226.25, the score he's trying to achieve.\[\Large\rm \frac{5}{8}\cdot 200+\frac{3}{8}x=226.25\]

zepdrix (zepdrix):

There's probably a more intuitive way to interpret it than that... But whatev :p So from there, you solve for x.

OpenStudy (anonymous):

270:)))))))))))))))))

OpenStudy (mathstudent55):

@zepdrix gave you correct solution to solve to find the answer. Here's how I think of it. Since 5/8 of the tournament is done, and he has 3/8 to go, let's assume he has played 5 games and will play 3 more.

OpenStudy (anonymous):

can you tell me if this is correct

OpenStudy (mathstudent55):

To average 200 in 5 games means he scored a total of x points which when divided by 5 gave an average of 200. x/5 = 200, or x = 1000

OpenStudy (mathstudent55):

You are correct.

zepdrix (zepdrix):

Yay good job lexi \c:/

OpenStudy (anonymous):

factor the sum or difference of two cubes 64p^3-1 my answer (4p+1)(16p^2-4p+1)

OpenStudy (mathstudent55):

He has 1000 points from the first 5 games. From the last three games he needs a total of points so that his overall average will be 226.25

zepdrix (zepdrix):

\[\Large\rm a^3-b^3=(a-b)(a^2+ab+b^2)\]Did you use the wrong formula maybe? :o

OpenStudy (mathstudent55):

For an 8-game average of 226.25, he needs 226.25 * 8 = 1810 total points. He already has 1000 from the first 5 games, so he needs 810 from the last 3 games. 810/3 = 270 The average for the last 3 games is 270.

OpenStudy (anonymous):

ok i getit more now thanks

OpenStudy (anonymous):

is the answer for other question (4p+1)(16p^2-4p+1)

OpenStudy (mathstudent55):